What the answer to (5t-1)-(4t+3)
Answer by lenny460(1073)

What the answer to (5t - 1)-(4t + 3)

(5t - 1) - (4t + 3)
-1 * 4t = -4t
-1 * 3 = -3
Now remove the parentheses:
5t - 1 - 4t - 3
Combine like terms:
5t - 4t = t
-1 - 3 = -4

Answer:
t - 4






Lennox Obuong
Algebra Tutor
Nairobi, Kenya
Email: obuong3@aol.com